OpenStudy (anonymous):

Solve the quadratic equation by using the square root property: (2x-4)^2 = 13

OpenStudy (anonymous):

(2x-4)(2x-4)= 13 4x2 - 8x - 8x +16= 13

OpenStudy (anonymous):

So now you just have to simplify.

OpenStudy (anonymous):

4x2-16x+3= y, I suppose you also want the roots of this equation?

OpenStudy (valpey):

\[(2x-4)^2=13\]Take square roots of both sides\[\sqrt{(2x-4)^2}=\pm\sqrt{13}\]\[(2x-4)=\pm\sqrt{13}\]Add 4 to both sides\[2x=4\pm\sqrt{13}\]Divide both sides by 2.

OpenStudy (anonymous):

Woops, hes right, I didnt use the square root property

OpenStudy (anonymous):

Valpey, I've tried entering in 2 +/- sqrt(13) several times, and the program says it's wrong. And yes, Ren, I do.

OpenStudy (anonymous):

Sorry...+/-sqrt(13/2)

OpenStudy (valpey):

@Ren is helping you convert this into the classic form of the quadratic equation. I don't think you need to do that. @Mbrecht Not sqrt(13/2) but sqrt(13)/2

OpenStudy (valpey):

\[x=2\pm\frac{\sqrt{13}}{2}\]

OpenStudy (anonymous):

No, I don't think I do. I seem to recall trying that as well, but I'll give it another spin.

OpenStudy (anonymous):

Are you entering both + and minus equations separately?

OpenStudy (anonymous):

No, I apparently hadn't...lol. Thank you both a tremendous amount.

OpenStudy (valpey):

Cheers.